### v1

# Learning Passive Continuous-Time Dynamics with Multistep Port-Hamiltonian Gaussian Processes

Abstract
We propose the multistep port-Hamiltonian Gaussian process (MS-PHS GP) to learn physically consistent continuous-time dynamics and a posterior over the Hamiltonian from noisy, irregularly-sampled trajectories. By placing a GP prior on the Hamiltonian surface $H$ and encoding variable-step multistep integrator constraints as finite linear functionals, MS-PHS GP enables closed-form conditioning of both the vector field and the Hamiltonian surface without latent states, while enforcing energy balance and passivity by design. We state a finite-sample vector-field bound that separates the estimation and variable-step discretization terms. Lastly, we demonstrate improved vector-field recovery and well-calibrated Hamiltonian uncertainty on mass-spring, Van der Pol, and Duffing benchmarks.
I Introduction
Learning physically consistent continuous-time dynamics is a core requirement for modeling and control, but noise and irregular sampling complicate the identification. Gaussian processes (GPs) provide a principled Bayesian framework for uncertainty quantification in system identification [1, 2], yet standard GP approaches typically ignore important structure such as energy conservation and passivity that are intrinsic to many mechanical and electrical systems. Recent work has begun to combine GP learning with port-Hamiltonian system (PHS) constraints, injecting physics-informed inductive biases via tailored kernels [3, 4]. These approaches directly enforce the port-Hamiltonian [5] structure in the learned dynamical model. However, these methods do not fully address the practical obstacles of measurement noise and irregular sampling in trajectories.
To close this gap, we introduce the multistep port-Hamiltonian system (MS-PHS) Gaussian process kernel. Our construction places a GP prior on the Hamiltonian surface $H(x)$ and leverages variable-step linear multistep methods (vLMM) to express discrete, irregularly-sampled trajectory constraints as finite linear functionals of $H$ and $\nabla H$. Since GPs are closed under linear operations, this yields closed-form posterior means and covariances for both the continuous vector field $f(x)=[J(x)-R(x)]\nabla H(x)$ and the Hamiltonian surface $H(x)$, directly from noisy, irregular data without leveraging a separate data preprocessing stage for extracting vector field information from trajectory data. By construction, the approach preserves PHS properties such as energy balance and passivity while accurately calibrating uncertainty from noisy, irregular trajectory observations.
We summarize a preliminary estimation-error decomposition that separates numerical truncation via the vLMM order and step size from GP interpolation uncertainty. The formal theorem and proof appear in a companion preprint [6], and are not required for the methods and experiments herein. Lastly, we empirically evaluate MS-PHS on three canonical oscillators (mass-spring, Van der Pol, Duffing) under noisy, irregular sampling, comparing against MS-ODE [7] and GP-PHS [3] (with Savitzky-Golay/LOESS filters). We report that across vector‐field recovery and Hamiltonian‐posterior calibration metrics, MS-PHS matches or exceeds the baselines while delivering uncertainty that tracks the true error, especially as noise and timestamp jitter increase.
I-A Related Works
The port-Hamiltonian system (PHS) formalism provides a unifying language for interconnection, energy balance, and passivity, with modular composition and stability guarantees [5, 8]. Building on this foundation, recent identification approaches learn models that retain passivity by design [3, 9, 10]. These advances clarify what physical invariants to encode when learning passive continuous-time systems, and naturally motivate probabilistic treatments under noisy, irregular sampling.
Gaussian processes (GPs) offer nonparametric priors with closed-form posteriors under linear functionals and have been used effectively for continuous-time vector-field recovery [11]. In many pipelines, derivative preprocessing and regular sampling are pragmatic choices that work well [3, 12, 13]. At the same time, when samples are irregular and end-to-end uncertainty quantification is essential, it is appealing to bring discretization inside the statistical model rather than treating it as a separate preprocessing stage.
The vLMM schemes such as AB/AM/BDF [14] translate flows into linear constraints in stacked vector field evaluations. Recent work shows that these constraints can be projected into GP posteriors for exact inference, even with variable step size [7]. This line of work substantially advances GP-based identification under irregular sampling. A key contribution of our work is to integrate physics-informed PHS priors [3] into the vLMM-based GP [7] in order to incorporate the noise and sampling irregularity into the posteriors of the Hamiltonian surface, enabling geometric uncertainty quantification with meaningful physical interpretations.
The paper is organized as follows: Section II reviews background on PHS and GPs. Section III develops the MS-PHS kernel and theoretical results. Section IV presents experiments.
I-B Notations
The notation $\mathbb{R}$ denotes the real number line. Vectors in $\mathbb{R}^{n}$ are column vectors. The kronecker product is denoted as $\otimes$. The notation $\rm vec$} denote the stacking operator that turns $X\in\mathbb{N}^{m\times n}$ into ${\rm vec}(X)\in\mathbb{R}^{mn}$. $A]_{ij}$ denotes the $i,j$ entry of a matrix $A$. We write ${\rm det}(\cdot)$ for the determinant of a matrix. Gradient of scalar function $H$ with respect to $x$ is denoted as $\nabla_{x}H$. We write $\mathbb{E}[\cdot]$ for expectation. The Gaussian process distribution is denoted as $\mathcal{GP}(\cdot,\cdot)$.
II Background
Necessary mathematical tools are introduced here.
II-A Port-Hamiltonian Systems
Energy-conserving and dissipative systems with input/output ports can be formally described as port-Hamiltonian systems (PHS) [5]:
$\begin{split}\dot{{x}}&=[J({x})-R({x})]\nabla_{{x}}H({x})+G({x}){u}\\ {y}_{\rm out}&=G({x})^{\top}\nabla_{{x}}H({x})\end{split}$
in which state ${x}(t)\in\mathbb{R}^{n}$ and the I/O ports ${u}(t),{y}_{\rm out}(t)\in\mathbb{R}^{m}$ evolve accordingly with time $t\in\mathbb{R}_{\geq 0}$. The Hamiltonian $H:\mathbb{R}^{n}\to\mathbb{R}$ is a smooth function that represents the total energy stored in the system. The interconnection matrix $J:\mathbb{R}\to\mathbb{R}^{n\times n}$ is skew-symmetric. The dissipative matrix $R:\mathbb{R}\to\mathbb{R}^{n\times n}$ is a positive semi-definite such that $R({x})=R^{\top}({x})\succeq 0$. The port mapping matrix $G:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ defines how an external input/output ports u{u}, ${y}_{\rm out}$ are coupled to the energy storage dynamics.
II-B Gaussian Process Regression (GPR)
A Gaussian process (GP) is a Bayesian prior over functions $f:\mathbb{R}^{d}\to\mathbb{R}$, where $f\sim\mathcal{GP}(m(\cdot),k(\cdot,\cdot))$, such that any finite collection of function values is jointly Gaussian [1]. The GP distribution, $\mathcal{GP}(m(\cdot),k(\cdot,\cdot))$, is fully specified by the mean and kernel function:
$m({x})=\mathbb{E}[f({x})],\quad k({x},{x}^{\prime})=\operatorname{Cov}[f({x}),f({x}^{\prime})].$ (1)
Given noisy observations $\mathcal{D}=\{({x}_{i},y_{i})\}_{i=1}^{N}$ with
$y_{i}=f({x}_{i})+\varepsilon_{i},\quad\varepsilon_{i}\sim\mathcal{N}(0,\sigma_{y}^{2}),$ (2)
the posterior over function values at test inputs $X_{*}=[{x}_{*}^{1},\dots,{x}_{*}^{L}]^{\top}$ is also Gaussian,
$f_{*}\mid\mathcal{D},X_{*}$ $\sim\mathcal{N}\bigl(\mu_{*},\Sigma_{*}\bigr),$ (3)
with posterior predictive mean and covariance:
$\begin{split}\mu_{*}&=m(X_{*})+K_{*X}\bigl(K_{XX}+\sigma_{y}^{2}I\bigr)^{-1}({y}-m(X)),\\ \Sigma_{*}&=K-K_{*X}\bigl(K_{XX}+\sigma_{y}^{2}I\bigr)^{-1}K_{X*},\end{split}$ (4)
where $X=[{x}_{1},\dots,{x}_{N}]^{\top}$, ${y}=[y_{1},\dots,y_{N}]^{\top}$, $K_{XX}]_{ij}=k({x}_{i},{x}_{j})$, $K_{*X}]_{ij}=k({x}_{*}^{i},{x}_{j})$, and $K]_{ij}=k({x}_{*}^{i},{x}_{*}^{j})$.
The noise variance $\sigma_{y}^{2}$ and kernel hyperparameters, e.g. length‑scales and output variance, are learned by minimizing the negative log marginal likelihood,
$\begin{split}-\log p({y}\mid X)&=\frac{1}{2}{y}^{\top}\bigl(K_{XX}+\sigma_{y}^{2}I\bigr)^{-1}{y}\\ &\qquad+\frac{1}{2}\log\det\bigl(K_{XX}+\sigma_{y}^{2}I\bigr)+\frac{N}{2}\log 2\pi.\end{split}$ (5)
For vector‑valued functions, such as continuous‑time dynamics ${f}:\mathbb{R}^{n}\to\mathbb{R}^{n}$, one can place independent GP priors on each output dimension or adapt a matrix kernel as detailed in [2].
II-B1 Port–Hamiltonian System Kernel
To enforce port-Hamiltonian dynamics on noisy vector field data $\{({x}_{k},{f}({x}_{k})+\varepsilon_{k})\}_{k=1}^{K}$, where:
${f}({x}_{k})\coloneq[J({x}_{k})-R({x}_{k})]\nabla_{{x}_{k}}H({x}_{k}).$
A zero‐mean GP prior is propose in [3] to place on the unknown Hamiltonian:
$H({x})\sim\mathcal{GP}\bigl(0,k_{\rm base}({x},{x}^{\prime})\bigr).$ (6)
Under the PHS dynamics (10), closure of GPs under affine transformations induces the matrix‐valued PHS kernel:
$k_{\rm phs}({x},{x}^{\prime})=\sigma_{f}^{2}J_{R}({x})\nabla_{{x}}\nabla_{{x}^{\prime}}k_{\rm base}({x},{x}^{\prime})J_{R}({x}^{\prime})^{\top},$ (7)
where $J_{R}({x})=J({x})-R({x})$ encodes the skew‐symmetric interconnection and dissipation matrices, and $\sigma_{f}^{2}\in\mathbb{R}_{>0}$ encode the signal noise. The above setup yields a GP prior:
$\dot{{x}}\sim\mathcal{GP}(G({x}){u},k_{\rm phs}({x},{x}^{\prime}))$ (8)
for the PHS dynamics in (10).
Remark II-B.1.
For concreteness, we use the squared exponential kernel for $k_{\rm base}$, while noting that other kernel choices are equally possible. Specifically,
$k_{\rm base}({x},{x}^{\prime})=\exp\Bigl(-\frac{1}{2}\sum_{i=1}^{n}\frac{({x}_{i}-{x}_{i}^{\prime})^{2}}{\ell_{i}^{2}}\Bigr),$ (9)
where $\ell=[\ell_{1},\dots,\ell_{n}]^{\top}$ is the vector of length-scales, with $\ell_{i}\in\mathbb{R}_{\geq 0},\ \forall i=1,\dots,n$. This corresponds to the automatic relevance determination (ARD) parameterization, which assigns an individual length-scale to each state dimension.
III Method
We formulate the problem and introduce our method here.
III-A Problem Formulation
Let $\mathcal{D}=\{(\tilde{{x}}(t_{k}),{u}(t_{k}),t_{k})\}_{k=1}^{K}$ denote the observations of the states and input signals of a PHS:
$\dot{{x}}=[J({x})-R({x})]\nabla_{{x}}H({x})+G({x}){u},$ (10)
sampled at time $t_{1},\dots,t_{K}$, where:
$\tilde{{x}}(t_{k})={x}(t_{k})+\varepsilon_{k},\quad\varepsilon_{k}\sim\mathcal{N}(0,\sigma_{{x}}^{2}I),$ (11)
with $\sigma_{{x}}^{2}\in\mathbb{R}_{>0}$ and x(t){x}(t) satisfying Equation (10) under initial condition ${x}(0)\in\mathbb{R}^{n}$. We follow the setting in [4] concerning our prior knowledge of (10):
  1. 1.
    the Hamiltonian $H\in\mathcal{C}^{\infty}$ is unknown;
  2. 2.
    the parametric structures of $J(\cdot),R(\cdot),G(\cdot)$ are known;
  3. 3.
    the parameters of $J(\cdot),R(\cdot),G(\cdot)$ are unknown.
We further assume that the PHS dynamics can be approximated using a specified numerical integration scheme [15].
III-B Multistep Port-Hamiltonian System Kernel
To perform exact Gaussian process inference of a Port–Hamiltonian vector field with irregularly-sampled trajectory data points, we combine the physics‐informed PHS prior from [3] with the variable step size multistep integrator  [7].
III-B1 Multivariate Variable Step Size Multistep Projection
Given noisy state observations $\{\tilde{{x}}_{k}\}_{k=1}^{K}$ at times $\{t_{k}\}_{k=1}^{K}$, an order-$(p)$ multistep integrator, e.g. Adams–Bashforth [14, Ch. 3.5], defines coefficient matrices, $A,B\in\mathbb{R}^{(K-M)\times K},$ such that the dynamic constraints of a control-affine system:
$\dot{x}(t)=f(x(t))+g(x(t))u(t),$
with $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$, ${g}:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ and $n=1$, can be enforced through:
$AX=B(\mathbf{f}(X)+\mathbf{g}(X)U),$ (12)
with the noiseless snapshot $X=[x_{1},\dots,x_{K}]^{\top}$ and $U=[{u}_{1},\dots,{u}_{K}]^{\top}$. The functions, $\mathbf{f}(X)=[f(x_{1}),\dots,f(x_{K})]^{\top}$ and $\mathbf{g}(X)U=[g(x_{1}){u}_{1},\dots,g(x_{K}){u}_{K}]^{\top},$ stacks the values of $f(x_{k})$ and $g(x_{k})u_{k}$, for $k=1\dots K$, accordingly. Notice that the first‐order forward Euler method with step size $h_{k}=t_{k+1}-t_{k}$ for zero-input systems is a special case of (12).
Example III.1 (Explicit Forward Euler).
The discretization:
$\dot{x}(t_{k})=f(x(t_{k}))\quad\Longrightarrow\quad x_{k+1}-x_{k}=h_{k}f(x_{k}),$
can be written in the form:
$AX=B\mathbf{f}(X),\quad\mathbf{f}(X)=\bigl[f(x_{1}),\dots,f(x_{K})\bigr]^{\top}.$
Here $A,B\in\mathbb{R}^{(K-1)\times K}$ have entries:
$A_{k,i}=\begin{cases}-1,&i=k,\\ +1,&i=k+1,\\ 0,&\text{otherwise},\end{cases}\quad B_{k,i}=\begin{cases}h_{k},&i=k,\\ 0,&\text{otherwise},\end{cases}$
so that: $(AX)_{k}=x_{k+1}^{\top}-x_{k}^{\top}$ and $(B\mathbf{f}(X))_{k}=h_{k}f(x_{k})^{\top}$.
Multistep integrators of order greater than or equal to $1$ (e.g., Adams-Bashforth/Moulton, BDF etc. [14, Ch. 3]) form updates from linear combinations of past states and vector-field evaluations. This class of numerical integrators yields the stacked linear constraint $AX=B\mathbf{f}(X)$ which preserves Gaussianity under a GP prior on $f$ and permits closed-form posterior updates [7] and recursive GP extensions [16].
When ${x}_{k}\in\mathbb{R}^{n}$ for $n\in\mathbb{N}$, the dynamic constraint encoded in (12) can be conveniently generalized to:
$A_{I}\coloneq A\otimes\mathsf{I}_{n},\ \$$B_{I}\coloneq B\otimes\mathsf{I}_{n},$ (13)
where $A_{I},B_{I}\in\mathbb{R}^{(K-M)n\times Kn}$, such that:
$Y\coloneq A_{I}{\rm vec}{(\tilde{X})}=B_{I}{\rm vec}{(\mathbf{f}(X)+\mathbf{g}(X)U)}+\boldsymbol{\varepsilon}.$ (14)
Remark III-B.1 (Unmodeled GP Input Noise).
In practice, the inputs $X$ appearing in $\mathbf{f}(X)$ and $\mathbf{g}(X)$ on the right-hand side of (14) are corrupted by measurement noise, so the GP is in fact heteroscedastic. A common remedy is to perform a local Taylor expansion of $f(x+\epsilon)$ around each datum and absorb the induced input noise into an input‐dependent output‐noise term, e.g. noisy input Gaussian process (NIGP), [17]. While higher‐order expansions and full posterior corrections [18, 19] can further improve accuracy, they lie outside the scope of this work. Accordingly, we assume noiseless inputs to the GP.
We now extend the exact GP inference framework proposed in [7] by injecting a physics-informed prior via the matrix-value PHS kernel and construct the multistep port-Hamiltonian system (MS-PHS) kernel:
$[K_{Y}]_{nm}$ $=\sum_{i=0}^{K-M}\sum_{j=0}^{K-M}[B_{I}]_{n,i}k_{\rm phs}({x}_{i},{x}_{j})[B_{I}]_{m,j},$ (15)
$[k_{Y}({x}_{*})]_{n}$ $=\sum_{j=0}^{K-M}[B_{I}]_{n,j}k_{\rm phs}({x}_{*},{{x}}_{j}),$ (16)
which is equivalent to the following matrix forms:
$K_{Y}=B_{I}K_{\rm phs}B_{I}^{\top},\qquad k_{Y}({x}_{*})=B_{I}k_{\rm phs}(X,{x}_{*}).$ (17)
The proposed MS-PHS kernel conditioning on $Y$ via the standard GP posterior formulas then yields closed‐form posterior means and covariances over the continuous‐time PHS dynamics, exactly accounting for both the physics prior and the multistep integration constraints.
MS-PHS inherits passivity and closure under power-conserving interconnection from GP-PHS [3, Prop. 1,2], a direct consequence of the closure of Gaussian processes under linear operators. Conditioning the GP prior on the multistep linear functionals yields a posterior GP with the same smoothness, so the gradients required for energy-rate and passivity calculations remain well-defined. Since the multistep projection acts componentwise and linearly on the port variables, it preserves energy additivity and the Port-Hamiltonian interconnection structure.
III-B2 Exact Inference via Multistep PHS Kernel
Given noisy observations $\mathcal{D}$, the posterior over the zero-input PHS vector field at a test input ${x}_{*}\in\mathbb{R}^{n}$ is:
${f}_{*}\mid\mathcal{D},{x}_{*}$ $\sim\mathcal{N}\bigl(\mu_{{f}},\Sigma_{{f}}\bigr),$ (18)
with posterior predictive mean and covariance computed as
$\mu_{{f}}$ $=k_{Y}({x}_{*})^{\top}\bigl(K_{Y}+\sigma_{{x}}^{2}A_{I}A_{I}^{\top}\bigr)^{-1}(Y-B_{I}G(X)U),$ (19)
$\Sigma_{{f}}$ $=k_{\rm phs}(x^{*},x^{*})-k_{Y}({x}_{*})^{\top}\bigl(K_{Y}+\sigma_{{x}}^{2}A_{I}A_{I}^{\top}\bigr)^{-1}k_{Y}({x}_{*}),$ (20)
where $K_{Y}=K_{Y}^{\top}\succ 0\in\mathbb{R}^{(K-M)n\times(K-M)n}$ is the training covariance from (15), $k_{Y}(x^{*})\in\mathbb{R}^{(K-M)n}$ is the cross-covariance from (16), $k_{\rm phs}(x^{*},x^{*})\in\mathbb{R}^{n\times n}$ is the raw PHS kernel evaluation at the test input, and $\sigma_{{x}}^{2}$ is the measurement‐noise variance defined in (11).
To analyze the error composition of MS-PHS GP, we recall two building blocks of the spectral-flow learning framework [6], vLMM theory and Spectral Regularization. An order-$p$ vLMM converts irregularly-sampled trajectory windows into linear constraints on stacked vector-field evaluations via a label map. Under the usual bounded step-ratio and order-$p$ consistency assumptions [14], the single-window local truncation error (LTE) for a vector field $f$, denoted as $\Delta_{\rm MS}(h_{k};f)$, satisfies the pointwise bound:
$\|\Delta_{\rm MS}(h_{k};f)\|_{2}\leq C_{\rm LTE}h_{k}^{p+1},$ (21)
with $C_{\rm LTE}>0$ independent of the step $h_{k}=t_{k+1}-t_{k}$.
Spectral regularization [20] on the windowed flow Hilbert space $\mathcal{W}$ supplies the statistical building block. Let $\mathsf{T}$ denote the population covariance operator on the windowed flow space and ${\mathsf{g}}_{\lambda}(\mathsf{T})$ a spectral filter with Lipschitz exponent $\mu$. Under a source condition of order $r>0$ and with $\ell$ labeled windows, standard perturbation and concentration arguments yield the finite-sample flow error rate, i.e., with probability at least $1-\eta$, [6, Thm. 1]:
$\|\widehat{\Phi}-\Phi_{\rho}\|_{\mathcal{W}}^{2}\leq C_{\rm flow}\log\frac{4}{\eta}\;\ell^{-\frac{2r}{2r+\beta}},\qquad\beta:=\max\{1,2\mu\}.$ (22)
where $\ell>0$ is the number of data points, $\widehat{\Phi},\Phi_{\rho}\in\mathcal{W}$ are the estimated and target flow of (10). Combining (22) with the vLMM LTE (21) produces a two-term high-probability vector-field bound:
Proposition 1 (Vector-field finite-sample high-probability bound).
Under the assumption that $H$ lies in the base RKHS induced by $k_{\rm base}$, for $0<\eta\leq 1$, the following inequality holds with probability at least $1-\eta$:
$\|\widehat{f}-f_{\rho}\|_{\mathcal{V}}^{2}\leq\frac{C_{\rm fit}}{c_{\rm obs}^{2}(h)}\log\frac{4}{\eta}\ell^{-\frac{1}{5}}+\frac{C_{\rm bias}}{c_{\rm obs}^{2}(h)}\ \mathbb{E}[h^{2p+2}],$ (23)
for some finite positive constant $c_{\rm obs}$, $C_{\rm fit}$, and $C_{\rm bias}$.
Proof.
Since Tikhonov regularization, e.g., GPR, has Lipschitz exponent $\mu=2$, we take $\beta=\max\{1,2\mu\}=4$. Under the assumption that $H$ lies in the base RKHS induced by $k_{\rm base}$, the baseline source condition $r=\frac{1}{2}$, i.e., $\Phi_{\rho}\in\mathrm{Range}({\mathsf{T}}^{\frac{1}{2}}))$ makes [6, Thm. 2] yield the rate exponent $2r/(2r+\beta)=\frac{1}{5}$. If stronger smoothness justifies $r=2$, the exponent improves to $2r/(2r+\beta)=\frac{1}{2}$ [20, Sec 3.2]. ∎
The estimated and target vector field are denoted as $\widehat{f},f_{\rho}\in\mathcal{V}$ respectively. The $(\ell,h)$-independent constants $C_{\rm fit},C_{\rm bias}$ are derived from $C_{\rm flow}$ and $C_{\rm LTE}$, respectively. The inverse term $1/c_{\rm obs}^{2}(h)$ is a caveat that arises from inverting the discretization methods when inferring the underlying vector field, which scales up with respect to decreasing step sizes $h$.
Proposition 1 shows that the estimation error splits into a statistical part that scales with $\ell$, and a discretization part that scales with $h$. Therefore, merely adding more data cannot remove discretization bias. Integrating the numerical integrator into the GP is necessary to obtain calibrated posteriors for vector fields and derived quantities like the Hamiltonian.
III-C Posterior Hamiltonian Surface
We now show how to recover the full posterior over the Hamiltonian surface $H(x)$, given our noisy observations $\mathcal{D}$ and a noiseless anchor $H(0)=H_{0}$. By placing a zero‐mean GP prior on $H$ with base kernel $k_{\rm base}(x,x^{\prime})$ in (6) and combining it with the VM‐PHS kernel for the trajectory data, we obtain a joint Gaussian prior over the augmented vector:
${y}_{\rm aug}=\begin{bmatrix}H(0)\\ {f}_{\rm vec}\end{bmatrix}\sim\mathcal{N}\bigl(0,K_{gg}\bigr),$
where ${f}_{\rm vec}\coloneq Y-B_{I}{\rm vec}(G(X)U)$ and the covariance matrix $K_{gg}$ is:
$K_{gg}=\begin{bmatrix}K_{HH}(0,0)&K_{Hf}(0,X)\\[6.0pt] K_{fH}(X,0)&K_{ff}(X,X)\end{bmatrix}$. (24)
Each block is constructed as follows:
  1. 1.
    Scalar block:
    $K_{HH}(0,0)=k_{\rm base}(0,0)+\epsilon_{H},$
    where $\epsilon_{H}$ is a small jitter ensuring positive definiteness.
  2. 2.
    Multistep gradient block:
    $K_{ff}(X,X)=K_{Y}+\sigma_{{x}}^{2}A_{I}A_{I}^{\top}.$
  3. 3.
    Cross‐covariance blocks:
    $K_{Hf}(0,X)=\begin{bmatrix}(J-R)(x_{1}),\nabla_{x_{j}}k_{\rm base}(0,x_{1})\\ \vdots\\ (J-R)(x_{1}),\nabla_{x_{K}}k_{\rm base}(0,x_{K})\end{bmatrix}B_{I}^{\top},$ (25)
    and $K_{fH}(X,0)=K_{Hf}(0,X)^{\top}$.
Once the hyperparameters, e.g., ARD lengthscales [1, 21], in $k_{\rm base}$ are learned through optimizing the negative log marginal likelihood (5), the surface posterior can be inferred by leveraging the linearity of differentiation operators [1, Ch. 9.4]. Given the a priori anchor $H(0)=H_{0}$ and gradient information ${f}_{\rm vec}$ encoded as the data transformed by the multistep projection, the posterior at any test point ${x}_{*}$ is obtained by standard Gaussian conditioning:
$H_{*}\mid\mathcal{D},H_{0},{x}_{*}\sim\mathcal{N}(\mu_{H},\sigma_{H}^{2}),$ (26)
where the posterior mean $\mu_{H}$ and variance $\sigma_{H}^{2}$ are:
$\mu_{H}({x}_{*})$ $=K_{*g}({x}_{*})\bigl(K_{gg}\bigr)^{-1}{y}_{\rm aug},$ (27)
$\sigma_{H}^{2}({x}_{*})$ $=k_{\rm base}({x}_{*},{x}_{*})-K_{*g}({x}_{*})\bigl(K_{gg}\bigr)^{-1}K_{*g}({x}_{*})^{\top},$ (28)
where:
$K_{*g}({x}_{*})=\bigl[k_{\rm base}({x}_{*},0),K_{Hf}({x}_{*},X)\bigr]\quad\in\mathbb{R}^{1+Kn}.$
In particular, $K_{Hf}({x}_{*},X)$ is constructed exactly as (25) but replacing the anchor with the test input ${x}_{*}$. Equations (27)–(28) then yield closed‐form expressions for the posterior mean and high‐probability variance of the Hamiltonian surface.
Remark III-C.1.
Since derivative information alone can recover the integrated surface only up to an arbitrary constant, anchoring $H(0)=H_{0}$ is essential to fix the additive constant of the GP. Any other noiseless constraint on $H$ (e.g., $H({x}_{i})=h_{i}$) can be incorporated analogously by augmenting $K_{gg}$ and ${y}_{\rm aug}$ with the corresponding rows and columns.
IV Experiments
We evaluate key features of the MS-PHS GP framework, including vector field and Hamiltonian surface posterior recovery. We learn the observation-noise variance and GP kernel hyperparameters by minimizing the negative log marginal likelihood (5) with ARD length-scales [1, 21], using ADAM optimizer [22]. All GP learning and quadratic programming modules are implemented using GPyTorch [23] and qpth [24] with PyTorch [25].
IV-A Benchmarks
Specifications of the three canonical oscillators as benchmarks for our posterior and recovery experiments: the Van der Pol oscillator, the Duffing oscillator, and a linear mass-spring-damper system are as follows:
Mass–Spring System
As a linear baseline, we choose stiffness $k=1.0$, mass $m=1.0$, and damping $d=0.0$ in:
$\ddot{q}=-\frac{k}{m}q-\frac{d}{m}\dot{q},$
representing a undamped harmonic oscillator.
Van der Pol Oscillator
We set the nonlinearity parameter to $\mu=1.0$ in the second‐order dynamics:
$\ddot{q}=\mu(1-q^{2})\dot{q}-q,$
which exhibits relaxation oscillations and non‐conservative behavior.
Duffing Oscillator
We use:
$\ddot{q}=-\alpha q-\beta q^{3}-\gamma\dot{q},$
where $\alpha=1.0,\ \beta=5.0,\ \gamma=0.5$, yielding a double‐well potential with nonlinear stiffness and damping.
To generate training and test data, each system is integrated over $t\in[0,20]$ using a classical fourth‐order Runge–Kutta solver with fixed time step $\Delta t=4\times 10^{-3}$. The resulting trajectories are then corrupted by additive state noise drawn from $\mathcal{N}(0,\sigma_{x}^{2}I)$ and irregularly subsampled at time points $t_{k}\in[0,20]$. For the irregularly subsampling scheme, we start from an $N$-point uniform grid $\{\tau_{k}\}_{k=1}^{N}$ on $t_{0},t_{1}]$ and add i.i.d. Gaussian “jitter” $\epsilon_{k}\sim\mathcal{N}\left(0,\sigma_{j}^{2}\right)$ with $\sigma_{j}=0.05$. The irregular timestamps are $t_{k}=\mathrm{clip}\left(\tau_{k}+\epsilon_{k},t_{0},t_{1}\right),$ after which they are sorted to enforce monotonicity. This produces near-uniform sampling with small, zero-mean perturbations, allowing mild clustering/gaps without changing the average sampling rate. Furthermore, we apply zero input ($u(t)\equiv 0$) on the Van der Pol oscillator and a nominal sinusoidal $\cos(\omega t)$ on the duffing oscillator and mass-spring system. We compare three inference schemes:
  1. 1.
    MS-PHS + Adams–Bashforth (MS-PHS-ab): our proposed multistep Port–Hamiltonian GP with Adams–Bashforth orders 1-3 [14, Ch. 3];
  2. 2.
    MS-ODE + Adams–Bashforth (MS-ODE-ab): the same multistep GP exact inference without a PHS prior with Adams–Bashforth orders 1-3, similar to the method proposed in [7];
  3. 3.
    GP-PHS + Savitzky–Golay (GP-PHS-savgol): a baseline that pre-processes the noisy data with a Savitzky–Golay filter [26] before fitting a PHS kernel [3]. We use Savitzky–Golay order 3 (savgol-3) in this paper.
  4. 4.
    GP-PHS + LOESS (GP-PHS-loess): a baseline that pre-processes the noisy data with locally estimated scatterplot smoothing (LOESS), a.k.a. locally weighted regression (LOWESS) [27, 28], before fitting a PHS kernel [3]. We use locally quadratic fitting (loess-2) in our simulations.
The Adams-Bashforth scheme is one of the simplest multistep integrators that fits the form in (12). Alternatives include Adam-Moulton [14, Ch. 3], backward-difference formulas [29, Ch. 5]. An in-depth comparison of these schemes for ODE-based exact GP inference appears in  [7].
We use LOESS with GP-PHS as a strong baseline for two reasons. First, the PHS kernel encodes a parametric Port-Hamiltonian structure with passivity/dissipation, injecting physics that provides a strong inductive bias for vector-field learning. Second, LOESS is a nonparametric, non-causal local polynomial smoother that naturally handles irregular sampling through distance-based kernel weights. Furthermore, LOESS is a distance-weighted local polynomial regression that works directly with irregular sampling grids. Savitzky-Golay can be seen as an evenly-weighted local polynomial variant of LOESS that works best with regular sampling grids.
The drawback of this family of preprocess filters is their look-ahead: LOESS is not directly compatible with online/recursive GP inference and real-time control. Moreover, separating the filter from GPR inhibits proper propagation of observational uncertainty into the posterior covariance. Below, we show that MS-PHS offers the best of both worlds, a physics-informed prior that reduces directional misalignment and a GP kernel that embeds the numerical integrator, enabling calibrated Hamiltonian uncertainty for downstream controls.
IV-B Vector Field Inference
In this section, we assess how well each method recovers the drift $f(x)$ from noisy, irregularly-sampled trajectories. Fig. 1 shows the performance measured by the mean-squared error (MSE) of the predicted vector field on an evaluation mesh around the simulated true trajectory. After $30$ runs per benchmark and method, we record the MSE and its 95% confidence interval across varying observation noise, time-grid jitter, and Gaussian process ARD initializations. In Fig. 1, we notice that the proposed MS-PHS GP regression with AB-3 integrator performs at least comparable to, if not better than, GP-PHS and MS-ODE in the three benchmarks. Furthermore, we observe that performance typically improves with increasing the order of the integrator.
Figure 1: Vector-field mean squared error (MSE) on three dynamical benchmarks—Mass-Spring, Van der Pol, and Duffing oscillators. Bars compare Gaussian-process regressors: multistep Port-Hamiltonian (MS-PHS-ab-1/2/3), multistep ODE (MS-ODE-ab-1/2/3), and a GP-PHS variant with Savitzky–Golay smoothing (GP-PHS-loess-2). Error bars indicate variability across runs. Lower is better.
To better understand the cause of error in Fig. 1, we visualize the learned flows against the ground-truth field on Van der Pol, overlaying posterior uncertainty to localize directional misalignment and data-sparsity effects in Fig. 2.
Figure 2: Van der Pol vector-field comparison. Ground truth (gray streamlines); MS-PHS GP (blue); MS-ODE GP (orange). Black dots are observations; red shading indicates high posterior standard deviation (std) of $f_{*}$, computed as $\sqrt{\sum_{i=1}^{n}(\Sigma_{{f}})_{ii}}$, with $\Sigma_{{f}}$ defined in (20). MS-PHS preserves the correct flow direction even in uncertain regions, whereas MS-ODE does not.
The added PHS structure: skew-symmetric interconnection and dissipative terms tied to $\nabla H$ constrain $f$ to be approximately tangent to the energy level sets while inducing damping. This physics prior regularizes extrapolation in data-sparse regions, reducing the spurious rotations seen with MS-ODE in Fig. 2.
To quantify directional agreement independent of scale, we next report the average cosine distance between true and learned vector fields with jittering grid effects turned off and replacing LOESS with Savitzky-Golay filter in GP-PHS. In Table I, each entry is the mean cosine distance, with the standard deviation in parentheses, computed over the evaluation mesh across runs.
TABLE I: Learned Vector Field Average Cosine Distances. Entries are “mean (standard deviation)” across 30 runs.
MS-PHS-ab-3 MS-ODE-ab-3 GP-PHS-savgol-3
MassSpring 0.001 (0.002) 0.001 (0.005) 0.003 (0.019)
VanDerPol 0.003 (0.018) 0.011 (0.034) 0.002 (0.016)
Duffing 0.003 (0.015) 0.042 (0.093) 0.003 (0.006)
Table I confirms the visual takeaway from Fig. 2, that physics-informed methods yield smaller cosine distances and tighter variability across runs. A short summary of Table I:
  1. 1.
    Mass–Spring: MS-PHS attains the best mean (tied with MS-ODE) but with a lower standard deviation, indicating more reliable alignment.
  2. 2.
    Van der Pol: Both PHS methods match on mean and clearly outperform MS-ODE on both mean and standard deviation.
  3. 3.
    Duffing: PHS methods are an order of magnitude better than MS-ODE, with low variance.
Overall, the added PHS structure reduces directional misalignment and stabilizes extrapolation in sparse regions, as anticipated.
IV-C Posterior Hamiltonian Recovery
We now move on to inspect the learned Hamiltonian geometry. Fig. 3 overlays the MS-PHS posterior mean and its $\pm 2\sigma_{H}$ band on the ground-truth Duffing surface. Uncertainty expands away from data (black points) and in high-curvature regions of the landscape, while remaining tight near visited states, consistent with the data support and the PHS prior.
Figure 3: 3D Hamiltonian landscape $H$ for the Duffing oscillator. The learned MS-PHS posterior mean $\mu_{H}$ surface (red) is overlaid with a reference/ground-truth surface $H_{\rm true}$ (green) and 96% posterior band $\mu_{H}\pm 2\sigma_{H}$ (translucent blue) for comparison. Black dots mark observed trajectory samples.
An important criterion for evaluating the Hamiltonian posterior is how well the learned Hamiltonian’s predictive variance tracks the true absolute error. Fig. 4 visualizes the true absolute error $|H_{\text{true}}-\mu_{H}|$ and the posterior standard deviation $\sigma_{H}$. Both quantities are small near the center of the explored region and grow toward the boundaries. Importantly, the 95% band $\mu_{H}\pm 2\sigma_{H}$ tracks the true error, indicating a well-calibrated posterior. We next compare the calibration of the Hamiltonian posterior variance for GP-PHS and MS-PHS.
Figure 4: Duffing oscillator Hamiltonian error and uncertainty over the $(q,\dot{q})$ plane. (Left) true absolute error between ground truth and learned mean, $|H_{\text{true}}-\mu_{H}|$. (Right) 95% posterior margin, $1.96\sigma_{H}$. The color bar shows magnitude; both error and uncertainty are small near the center and increase toward the domain boundary. Comparing the left panel and right panel shows that the $95\%$-posterior credible band ($\mu_{H}\pm 2\sigma_{H}$) tracks the true error: $|H_{\text{true}}-\mu_{H}|$.
Figure 5: Scaling of Hamiltonian posterior error and uncertainty with observation noise on the Duffing system. For each noise level $\sigma_{x}$, points show the median and bars the interquartile range (log scale). Panels share the y-axis for direct comparison. (Left) MS-PHS-ab-3: $\sigma_{H}^{2}$ rises in step with $H_{\text{MSE}}$, indicating calibrated uncertainty across noise. (Right) GP-PHS-loess-2: $H_{\text{MSE}}$ increases sharply while $\sigma_{H}^{2}$ lags, especially for $\sigma_{x}^{2}\geq 0.02$, suggesting underestimation of uncertainty.
Figure 6: Error-uncertainty Ratio of Hamiltonian uncertainty on Duffing. Boxplots show the distribution of the Error-uncertainty ratio $H_{\rm mse}/\sigma^{2}_{H}$ across observation noise levels $\sigma_{x}$, where dashed line represents ideal calibration at 1. MS-PHS-ab-3 remains near or below 1 across $\sigma_{x}$, indicating posterior variance that tracks the error, while GP-PHS-loess-2 drops below 1 as noise increases, evidence of overconfident posterior estimates.
In Fig. 5, GP-PHS-savgol-3 shows error $H_{\mathrm{MSE}}$ increasing rapidly with noise, whereas its reported uncertainty, summarized by $\sigma_{H}^{2}$, lags, especially for $\sigma_{x}^{2}\geq 0.02$. This mismatch is consistent with the decoupling between the LOESS smoothing stage and the GP kernel: the pre-processing can absorb and bias modeling errors that the GP posterior never sees, so the predictive variance fails to reflect the true model error. By contrast, MS-PHS-ab-3 exhibits posterior variance that accurately tracks the realized error. In Fig. 5 (left), $\sigma_{H}^{2}$ rises in step with $H_{\text{MSE}}$ across noise levels, indicating that the posterior dispersion scales with the actual reconstruction difficulty.
To further investigate the issue of calibrating posterior variance to match MSE of Hamiltonian, we leverage the notion of error-uncertainty ratio, defined as $H_{\rm mse}/\sigma^{2}_{H}$. The posterior with $H_{\rm mse}/\sigma^{2}_{H}<1$ is conservative, $H_{\rm mse}/\sigma^{2}_{H}>1$ is overconfident, $H_{\rm mse}/\sigma^{2}_{H}=1$ is well calibrated. The calibration ratio in Fig. 6 makes this explicit: the values concentrate near, or slightly below, the ideal line at $1$ for MS-PHS-ab-3; while GP-PHS-savgol-3 rise above $1$ as noise increases, i.e., uncertainty is underestimated.
TABLE II: Ratio of Hamiltonian posterior variance to MSE ($\sigma_{H}^{2}/H_{\mathrm{MSE}}$) across time-grid jitter levels on Duffing. Entries are “median [Q1, Q3]” across 30 runs.
$\sigma_{j}$ MS-PHS-ab-3 GP-PHS-loess-2
0.00 1.16 [0.847, 1.42] 4.24 [3.28, 5.33]
0.01 1.07 [0.831, 1.29] 10.1 [8.96, 11.4]
0.02 1.14 [0.929, 1.39] 9.06 [7.54, 12.2]
0.05 1.04 [0.903, 1.36] 9.52 [7.75, 10.5]
0.10 1.06 [0.927, 1.38] 9.2 [7.22, 11]
We observe similar phenomena under time-grid jitter. Table II reports the ratio $\sigma_{H}^{2}/H_{\mathrm{MSE}}$ across jitter levels $\sigma_{j}^{2}$. MS-PHS-ab-3 stays close to unity, medians $\approx 1$ with relatively tight interquartile ranges, again signaling calibrated variance. In contrast, GP-PHS-savgol-3 remains far below $1$ for all jitters, confirming that the separation between smoothing and inference leaves its predictive variance insensitive to the induced modeling error.
V Conclusion
We introduced the multistep Port-Hamiltonian Gaussian process (MS-PHS GP) kernel for learning continuous-time dynamics from noisy, irregularly-sampled trajectory data while enforcing port-Hamiltonian structure by construction. By placing a GP prior on the Hamiltonian and embedding variable-step multistep integrator constraints as finite linear functionals, MS-PHS yields closed-form posterior means and variances for both the vector field and the Hamiltonian surface. Theoretically, the construction preserves passivity and compositionality under power-conserving interconnection, and the estimation error decomposes into a multistep truncation term (controlled by integrator order and step size) plus a GP interpolation term (controlled by posterior uncertainty).
Empirically, MS-PHS (particularly with higher-order Adams–Bashforth projectors) recovers vector fields and Hamiltonian geometry more accurately and with better directional alignment than non-physics or prefiltered baselines, while producing well-calibrated posterior uncertainty that tracks empirical error such as $H_{\rm mse}$ across noise and sampling jitter. These properties make MS-PHS especially attractive for downstream control tasks where physically consistent models and reliable uncertainty quantification are essential.